### v2

# ComplexTempQA: A Large-Scale Dataset for Complex Temporal Question Answering

Abstract
We introduce ComplexTempQA, [^1] Dataset and code are freely available at: https://github.com/DataScienceUIBK/ComplexTempQA a large-scale dataset consisting of over 100 million question-answer pairs designed to tackle the challenges in temporal question answering. ComplexTempQA significantly surpasses existing benchmarks like HOTPOTQA, TORQUE, and TEQUILA in scale and scope. Utilizing data from Wikipedia and Wikidata, the dataset covers questions spanning over two decades and offers an unmatched breadth of topics. We introduce a unique taxonomy that categorizes questions as attributes, comparisons, and counting questions, each revolving around events, entities, and time periods. One standout feature of ComplexTempQA is the high complexity of its questions, which demand effective capabilities for answering such as across-time comparison, temporal aggregation, and multi-hop reasoning involving temporal event ordering and entity recognition. Additionally, each question is accompanied by detailed metadata, including specific time scopes, allowing for comprehensive evaluation and enhancement of the temporal reasoning abilities of large language models. ComplexTempQA serves both as a testing ground for developing sophisticated AI models and as a foundation for advancing research in question answering, information retrieval, and language understanding.
1 Introduction
Figure 1: Example types of temporal reasoning in questions sampled from ComplexTempQA (left) with the required inference steps (right) and timeline-based visualization (above).
Temporal Question Answering (TQA) is about answering questions that require an understanding of and reasoning about temporal knowledge [18, 22, 21, 30]. This makes it substantially different from traditional Question Answering (QA) (see Figure 1). The complexity of questions users tend to ask to search systems nowadays has increased. Unlike traditional keyword-based searches, directly posed questions  [16] of increasing complexity are now essential for evaluating QA as well as TQA approaches, driven by the advancements in the current-generation LLMs.
The few existing datasets for TQA, such as TORQUE [30], ArchivalQA [47], ChroniclingAmericaQA [33] and TEQUILA [21] fall however short in several regards: First, these datasets are limited in size, typically containing only a few thousand questions. This poses significant challenges for effectively training LLMs, as illustrated in  [30, 20, 27, 31, 49, 29]. Second, while these datasets predominantly focus on specific question categories about entities or time periods, they lack comprehensive coverage across a wide variety of question types. Third, existing datasets generally include only straightforward questions and miss complex inquiries that necessitate unfolding and inference to generate accurate responses, thereby limiting trained models’ ability to handle sophisticated temporal reasoning tasks [30, 31, 49, 29]. Fourth, the datasets generally cover only short periods of time and lack features such as difficulty ratings for questions and the ability to filter data by specific time periods, which is essential for detailed and customized temporal studies.
To address the identified challenges, we introduce ComplexTempQA, a novel dataset that surpasses existing resources in scale and complexity, designed specifically for the analysis, training, and testing of LLMs and QA systems on complex temporal knowledge: (1) Size: ComplexTempQA comprises over 100 million question-answer pairs, making it by far the largest dataset in temporal QA. (2) Question Types and Taxonomy: Our dataset includes diverse categories of questions such as attribute-, comparison-, and counting-type questions, each relating to events, entities, or time periods. We generate questions at scale based on facts from Wikidata and typical question types identified from Wikipedia, ensuring a broad coverage of both domains and question types. In addition, we assure a broad thematic scope by employing the IPTC Media Topics taxonomy to further diversify topics. (3) Complexity and Timeframe: Questions in ComplexTempQA require advanced reasoning skills, including event/entity matching, multi-hop question answering, across-time comparison, and temporal ordering. We set the time frame from 1987 to 2023 which is sufficiently long for diachronic studies and exploration of continual learning approaches for ML architectures and LLMs. (4) Metadata and Evaluation: Each question is provided with detailed metadata, such as the time period it covers and a difficulty rating that classifies questions as “easy” or “hard” depending on the question type and expected familiarity of the entities involved. This metadata enables an accurate assessment of the ability of language models in coping with the change of time focus.
ComplexTempQA serves multiple purposes in advancing the study and application of LLMs in temporal reasoning. Primarily, it enhances the analysis of LLMs’ performance over time, addressing the critical need to understand temporal factual knowledge, identify temporal blind spots, and evaluate temporal commonsense reasoning capabilities [45, 50]. Beyond performance evaluation, this dataset acts as a foundational platform for developing further question-generation tools. Our structured taxonomy aids in refining and developing taxonomies for specialized tasks and datasets. For targeted research, subsets of ComplexTempQA can be used, applying specific filters to focus on particular aspects of temporal reasoning and specific time frames.
A crucial aspect of our work involves assessing how state-of-the-art LLMs handle the complex temporal questions presented by ComplexTempQA. In Section 6, we evaluate a range of models, from those with 7 billion to 175 billion parameters, using methods including zero shot, few shot, and retrieval-augmented generation (RAG). The evaluations provide insights into the current capabilities and limitations of LLMs in processing temporal information.
Overall, we make the following contributions in this paper:
  1. 1.
    We present and release ComplexTempQA,11footnotemark: 1 a large-scale dataset with over 100 million question-answer pairs for temporal question-answering based on a new taxonomy of temporal question types.
  2. 2.
    We detail our methodology for dataset creation, involving data retrieval from Wikidata, event extraction from Wikipedia, complexity assurance and rating, as well as temporal ambiguity based filtering. The dataset can be also easily extended to encompass other time periods.
  3. 3.
    We evaluate diverse LLMs on our dataset using zero shot, few shot prompting as well as RAG approaches, providing insights into their performance in handling temporal questions.
2 Related Work
Table 1: Comparison of ComplexTempQA with existing datasets highlighting the key aspects of question creation methodologies, answer types, complexity, temporal scope, and structure. Attr means attribute-type question, Comp denotes comparison-type questions and Count are counting-type questions.
Dataset Temporal Questions #Questions Creation Method Source Answer Type Complex Question Types Question Type Time Frame Temporal Metadata Multi-Hop SQuAD 2.0 [34] No 158K Crowd sourced Wikipedia Extractive - Attr, Comp Unspecified No No CNN/Daily Mail [7] No 1,385K Aut. Generated News Cloze - Attr Unspecified (short) No No NewsQuizQA [25] No 20K Crowd sourced News Multiple choice - Attr 2018 - 2022 No No NewsQA [42] Partially 119K Crowd sourced News Extractive TimeAttr Attr 2007-2015 No No HOTPOTQA [52] No 113K Crowd sourced Wikipedia Extractive - Attr, Comp Unspecified No 2 hops LC-QuAD 2.0 [11] Partially 30K Crowd sourced Wikipedia Extractive TimeAttr Attr, Count Unspecified No No TORQUE [30] Yes 21K Crowd sourced News Generative TimeAttr Attr Unspecified (short) No No Time-Sensitive-QA [8] Yes 41K Aut. Generated Wikipedia Extractive TimeAttr Attr Unspecified (long) No No TempQuestions [20] Yes 1K Aut. Generated Freebase Extractive TimeAttr, TimeComp, TimeCount Attr, Comp, Count History No No TKGQA [31] Yes 5K Aut. Generated News Extractive TimeAttr Attr 2022 No No MenatQA [49] Yes 2K Aut. Generated Wikipedia Extractive TimeAttr Attr Unspecified (long) No No TDDiscourse [29] Yes 6K Aut. Generated News Extractive TimeAttr Attr Unspecified (short) No No ArchivalQA [47] Partially 532K Aut. Generated News Extractive TimeAttr Attr, Count 1987-2007 No No ComplexTempQA Yes 100,228K Aut. Generated Wikipedia Extractive, Boolean TimeAttr, TimeComp, TimeCount, Unnamed questions Attr, Comp, Count 1987-2023 Yes $\leq$ 2 hops
Table 1 gives an overview of question answering datasets including ones applicable for TQA (indicated by the "Temporal Questions" column) and outlines their differences compared to the ComplexTempQA dataset. The table reveals a notable discrepancy in question volume compared to our dataset, which substantially surpasses others. The largest datasets predominantly rely on automated creation methods, drawing from news articles or Wikipedia rather than crowd-sourcing. Upon closer examination, we find that while several datasets contain complex temporal questions, they largely revolve around TimeAttr inquiries, focused on relatively simple questioning about the duration and order of events or entities. Furthermore, only HOTPOTQA [52] and ours incorporate question hops distinguishing them from other datasets. However, HOTPOTQA does not contain temporal questions.
TORQUE [30] introduces questions emphasizing temporal relations such as "before," "after," and "start". The authors trained a model to evaluate questions specifically based on these temporal constraints. QA TempEval dataset [27] has been designed with the focus on temporal entities and temporal relations, which are easier to generate automatically. TEQUILA [21] uses temporal expressions like dates or implicit temporal signals such as "before" or "after." In addition to these works, TempQuestions [22] has been released as a temporal questions benchmark, containing 1,271 questions that are all temporal in nature, paired with their answers. That work provides a simple definition for temporal questions and demonstrated the need for further research on complex questions. To improve question answering for event-focused questions in temporal collections of news articles, Wang et al. [46] propose a large-scale question-answering model designed specifically for long-term news article collections. The authors focus mainly on introducing an effective approach for re-ranking articles by using document publication dates and temporal expressions embedded in text. Stricker [38] leverage a popular approach used for general question answering and answer extraction to find answers to temporal questions within paragraphs.
Several researches have leveraged knowledge graphs to generate questions and find accurate answers. LC-QuAD 2.0 [11] demonstrated the potential for automating the generation task without human involvement by querying the Wikidata knowledge graph via SPARQL. A similar strategy was employed by Cui et al. [10], who used Wikidata’s SPARQL endpoint to create question-answer pairs based on questions from Compositional Freebase Questions. For our work, we chose Wikidata as our endpoint due to its extensive database covering a wide range of topics [37, 34, 39].
Unlike other datasets, ours stands out due to its unique characteristics: (a) it comprises a several magnitudes larger number of question-answer pairs, (b) it categorizes these questions into specific types, (c) it integrates complex questions, and (d) all questions have a temporal nature, with a strict assignment to a specific time span.
3 Dataset Characteristics
In this section, we describe the ComplexTempQA characteristics along the four dimensions.
(1) Size:  ComplexTempQA comprises 100m question-answer pairs and covers the period from 1987 to 2023. The dataset has been curated to probe the understanding of temporal knowledge within this 36-year span, encapsulating events, entity milestones, and other time-sensitive information.
(2) Question Types and Taxonomy: The primary objective when constructing ComplexTempQA was to incorporate a broad spectrum of temporal questions. Our dataset offers extensive coverage across diverse subjects by aligning with IPTC MediaTopic standards, ensuring a broad range of relevant topics. This adherence guarantees both temporal comprehensiveness and topical diversity. The questions are organized using a taxonomy that categorizes them by the questioned entity or event and the nature of the knowledge they probe, divided into attribute, comparison, and counting questions. The numbers for each question type are shown in Table 3. Below we provide description of each type, while specific examples are given in Table 4 in Appendix A.
Figure 2: Dataset distribution (time questions are integrated within events or entities).
Name Total Attribute Event 83,798 Attribute Entity 84,079 Attribute Time 9,454 Comparison Event 25,353,340 Comparison Entity 74,678,117 Comparison Time 54,022,952 Counting Event 18,325 Counting Entity 10,798 Counting Time 12,732 Multi-Hop: 76,933 Unnamed Event: 8,707,123 Total: 100,228,457
Figure 3: Taxonomy of temporal question types.
Attribute Questions. Attribute-type questions pertain to the attributes of events or entities, or are related to a specific time (e.g., “When was the fall of the Berlin Wall?”). The answers to these questions typically consist of events, entities, dates (as well as their lists) or real numbers. To ensure uniformity in the interpretation of numerical values across responses, we convert all units into the International System of Units (SI units) for questions that solicit a numerical answer involving units.
Comparison Questions. Comparison-type questions involve the comparison of one or two events, entities, or time periods. The comparison aspect can be either numerical or temporal. For example, we compare two temporal attributes in the question: “Did Halabja chemical attack happen after John F. Kennedy Jr. plane crash?”. The answer to these questions can be either true, false, an entity, or an event.
Counting Questions. Counting-type questions - a kind of aggregation questions - record the frequency of a particular event or the frequency of an attribute of an entity appearing. An example is “How often did an aviation accident with more than 38 deaths between 1987 and 2023 in Thailand happened?”. To make the answering of the counting-type questions more precise, we always consider a specific time period and in some cases an attribute’s threshold. The threshold is determined by calculating the average number across all data in the Wikidata pertaining to the specific item. For instance, in the example provided, the threshold pertains to the number of deaths associated with the event. The answer to these questions is always a natural number.
(3) Difficulty of Questions.  We incorporate questions of varying difficulty levels, where complex questions are typically derived from multi-hop queries or are modified versions of standard questions that implicitly reference the event without stating its name. We call the latter, unnamed event questions (see Sec. 4.2 for details). All the questions in ComplexTempQA are categorized into two levels: easy and hard. As mentioned above, difficulty is determined either by the number of page views of the Wikipedia page related to the events or entities, or by the inherent complexity of the question. Counting-type questions and other complex queries such as multi-hop or unnamed questions are automatically classified as hard. For all other question types, difficulty is assessed using page view statistics from the English Wikipedia. Events or entities with higher visibility, evidenced by more page views, are generally categorized as easy, reflecting their greater familiarity and accessibility to a wider audience. To compute the difficulty rating scores, a question is only categorized as easy if all of its containing entities or events are rated as easy. An example of an easy time attribute-type question is "When was the death of Diana, Princess of Wales?", while an example of a hard time attribute-type question is "When was the Purley station rail crash?".
(4) Metadata and Evaluation.  ComplexTempQA includes a set of basic as well as, when available, several additional metadata fields for each question. The metadata serves multiple purposes, including facilitating the retrieval of entities and events from the question or answer. Additionally, it enables in-depth analysis of the entire dataset and facilitates easier segmentation based on attributes such as type, year or difficulty rating.
Firstly, in the metadata, we specify the type of question based on the taxonomy shown in Figure 3. Next, we include the identifier numbers of all corresponding Wikidata items that are either questioned or are part of the answer. Furthermore, we add the country, if applicable, for both the questioned entities or events and for their answers. We also include complexity indicators for the questions, such as the properties used for creating hop questions and whether it is an unnamed question or not. Additionally, we rate the difficulty of the question, indicating whether it is considered difficult or easy. Lastly, we include a time span to specify the temporal range of the question. In Appendix B we provide more information on each metadata as well as we showcase an example.
4 Dataset Creation Pipeline
For constructing the dataset, Wikipedia served as the primary source for event attributes and comparison-type questions, while Wikidata provided the source for the entity attributes and counting-type questions. In the following, we provide a detailed description of the dataset construction process. In Appendix D we also present a detailed view of the dataset creation pipeline.
4.1 Dataset Source Extraction.
As the source of the event attribute-type questions, we extracted every entry from single year pages from Wikipedia (see example [^2] E.g., https://en.wikipedia.org/wiki/1987) to gather information on significant events that happened over a span of 36 years (1987 to 2023). As for entity attribute-type questions, we compiled a comprehensive list of entities by querying Wikidata using SPARQL, focusing on items such as “movies” and “heads of state”. Those predefined items are derived from the IPTC MediaTopic standards [^3] https://show.newscodes.org/index.html?newscodes=medtop&lang=en-GB&startTo=Show which have been manually created. Since there is a large number of entries for each item, we further refined the results based on page views. For instance, over 100,000 entities were extracted for “movie”, which we further refined into 35,991 entities based on their page views from their corresponding Wikipedia pages.
For counting-type questions, a different approach was necessary because we do not operate on single events or entities anymore, but we count events or entities, as follows. We first create a SPARQL query templates tailored to individual counting types using Wikidata. For instance, by querying the Wikidata item "Presidents of the United States" within a specified year range, we obtained a list of all Presidents of the United States recorded within that time frame as the potential answer for a counting-type question. In some cases, like the previous example, we could directly use such a list. However, in other cases, the list might be incomplete due to the omission of minor events not documented in Wikidata, such as "earthquakes." Therefore, for the counting-type questions, we introduced a threshold. This threshold attribute, such as the "number of deaths of some calamity", is derived based on the average values of corresponding attributes computed from the Wikidata.
4.2 Dataset Complexity Enhancement.
In order to increase the complexity of the questions, we introduce a module for creating multi-hop questions, by leveraging common attributes of entities and events extracted in the previous step. Concretely, we created multi-hop questions by combining up to two attributes of events and entities to form a question structure with a maximum of two hops.
To introduce further complexity, we created additional events based on those we already gathered. In particular, we generated new expression about the same event, but with implicit naming. For instance, the event "Lauda Air Flight 004 accident" was rephrased as "the aviation accident in 1991 which resulted in 223 deaths." We used a property and the year of the event to create the new unnamed version. To ensure the uniqueness of the new unnamed entry and avoid referencing other events due to potential ambiguity, we executed a subsequent SPARQL query to verify that no other entities or events match the specified attributes, allowing us to build an unambiguous catalog of events and entities.
4.3 Dataset Formation and Enrichment.
The next step involves the creation of the actual dataset of question-answer (QA) pairs by integrating the events, entities, and counting results into predefined question templates (see Appendix A), described as follows.
  • attribute-type questions: we queried specific attributes of an entity or event, enhancing the attribute sets with additional relations as introduced in [8]. The addition of these relations is useful for enhancing the comprehensiveness of the queries.
  • comparison-type questions: we selected up to three events or entities to compare, employing terms such as “higher/lower”, “before/after”, or “happened first” to frame the queries.
  • counting-type questions: we specified a time frame (e.g., 5 years) and further inquired the categories (e.g., earthquakes, Nobel Peace Prize) associated with the events and entities.
Once the QA pairs are formed, we further enriched them with additional metadata including the corresponding Wikidata IDs, the relevant country, and the specific time frame of each question. Additionally, we assigned difficulty ratings as outlined in Sec. 3. This rating system, combined with the metadata from the entities and events, adds an extra layer of informational depth to the dataset.
Figure 4: Result of the quality check.
Figure 5: Performance of Zero shot LLMs.
Method Parameters Precision Recall F1 Con Zephyr 7B 3.76 33.50 4.90 55.97 Falcon 7B 0.31 34.22 0.62 64.21 Llama-chat 7B 7B 3.68 33.94 6.09 50.32 Mistral 7B 3.73 48.34 6.33 58.31 LLama-chat 13B 13B 3.61 32.77 6.05 53.27 Vicuna 33B 2.02 37.27 3.63 58.26 Mixtral 8x7B 3.34 40.41 5.65 60.21 LLama-chat 70B 70B 5.19 39.30 8.31 57.33 Wizardlm 70b 1.63 27.63 2.80 50.44 GPT-3.5 175B 4.71 34.46 7.74 49.31
5 Quality Check
A quality check by human raters is essential to ensure the quality of ComplexTempQA. This process verifies that the dataset meets predefined criteria for clarity and readability. By leveraging human expertise, we can identify and address potential errors, ambiguities, or biases, thereby enhancing the dataset’s reliability and trustworthiness.
We perform a user study, in which we asked 11 volunteers (4 females and 7 males, with 4 having secondary education and 7 possessing postgraduate degrees; ages ranging from 26 to 56 years old, predominantly within their 20s or 30s) to assess 450 randomly selected questions, which were evenly distributed across all question types. The participants were asked to evaluate those questions according to four metrics as follows: (1) the readability of the question, (2) how easy it is to answer the question before using a web search engine, (3) how easy it is after conducting a web search, and (4) the overall clarity of the question including ambiguity. The participants had to assign a rating on a scale of 1 to 5 (Likert scale). For (2) and (3), the participants were first asked to rate the ease of answering the questions based solely on their own knowledge. Subsequently, they provided a rating on how easy it was to find the answer after using a search engine. Additionally, participants were explicitly instructed not to utilize any LLMs when providing their answers.
The results are given in Figure 5. The evaluation revealed that participants rated the readability and clarity considerably high. We also observed that answering questions without resorting to a search engine was challenging. After conducting web searches, participants on average found answering questions considerably easier. Counting-type questions remained most difficult to answer, even after searching online since often there are no immediately available answers for such questions but rather they must be inferred by conducting some online research. Clarity was a bit low-rated for some counting- and comparison-type questions, suggesting a need for reformulation in future iterations. Finally, the rating system was found to significantly impact the scores with hard questions being more difficult to answer, both before and after conducting web search.
6 Evaluation
We evaluate the performance of multiple LLM models, with a particular focus on the taxonomy of our questions to gain deeper insights into how the models fare across different question types. We have selected an equal number of samples from each question type, ensuring a balanced representation of types and complexities. The total number of questions we used was 12,0001200012,00012 , 000. The prompts utilized for the models are detailed in Appendix C.
LLM Models.  We evaluate various state-of-the-art LLM models: (1) Llama is optimized for dialogues and employs supervised fine-tuning and reinforcement learning for alignment with human preferences [41]. (2) Mistral is known for its power efficiency, using instruction fine-tuning to enhance performance on multiple benchmarks [23]. (3) Mixtral utilizes a Sparse Mixture of Experts approach, offering superior inference throughput by leveraging up to 45B parameters, but only about 12B during inference [23]. (4) GPT3.5, an iteration of GPT3 with 175B parameters, designed to reduce toxic output generation and trained on a diverse corpus up until December 2023 [4]. (5) Falcon focuses on efficiency with modern architecture (e.g., multi-query attention), trained on over 1T tokens of the RefinedWeb corpus [32]. (6) Vicuna is a conversational model fine-tuned on user-shared conversations, using an autoregressive, decoder-only architecture [28]. (7) Zephyr employs distilled supervised fine-tuning for improved accuracy, trained as part of the Mistral lineup [43]. (8) Wizardlm has been fine-tuned for following complex instructions [51].
Evaluation Metrics.  Evaluating LLMs, especially for question answering, is challenging due to the verbose nature of responses. Traditional metrics like Exact Match and F1 score may not be suitable when used alone. To address this and to follow related works [14, 3], we use also model-agnostic metrics like Token Recall and Answer String Containment [1]. Token Recall measures how well the model’s response covers the ground truth. Answer String Containment assesses if the model’s response contains the core answer.
We evaluate the LLM’s capabilities regarding (a) zero shot QA, (b) few shot QA, (c) RAG QA, and (d) the effectiveness of LLMs in different QA settings. Furthermore, we divide the performance metrics by question type to gain deeper insights into how each type performs individually.
Figure 6: Performance of Few shot LLM Models.
Method Parameters Shots Precision Recall F1 Con EM Llama-chat 7B 0 3.675 33.935 6.09 50.315 0.035 1 7.08 27.21 9.00 43.09 3.57 2 23.05 30.65 23.655 33.06 21.49 3 23.67 30.38 24.22 31.78 22.25 Llama-chat 13B 0 3.605 32.76 6.05 53.27 0.0085 1 22.865 27.91 23.345 38.11 21.525 2 31.645 32.35 31.56 31.465 30.37 3 30.37 32.03 30.495 32.46 29.00 Llama-chat 70B 0 5.191 39.30 8.31 57.32 0.138 1 25.74 32.61 18.8 49.60 13.61 2 34.01 44.21 35.26 43.27 31.78 3 37.09 46.57 38.44 42.77 34.59 Mistral-Instruct 7B 0 3.73 48.33 6.325 58.30 0.034 1 24.68 35.33 25.87 41.81 21.83 2 32.74 35.91 33.145 34.82 30.55 3 35.28 37.91 35.68 36.26 32.93 Mixtral 8x7B 0 3.34 40.41 5.65 60.20 0.156 1 6.76 39.06 9.17 52.87 2.75 2 14.27 41.89 16.44 52.64 10.22 3 15.49 44.03 17.83 51.57 11.25
Figure 7: Performance of LLMs in RAG QA setting.
Method Parameters Context Precision Recall F1 Con Llama-chat 7B No Context 3.67 33.93 6.09 50.31 Retriever 3.59 33.67 5.97 53.48 True Context 3.92 37.40 6.49 56.03 Llama-chat 13B No Context 3.60 32.76 6.05 53.27 Retriever 3.50 34.22 5.84 55.38 True Context 3.75 37.09 6.28 57.42 Llama-chat 70B No Context 5.19 39.30 8.31 57.32 Retriever 5.27 36.16 8.12 56.45 True Context 5.82 38.59 8.82 58.26 Mistral-Instruct 7B No Context 3.73 48.33 6.32 58.30 Retriever 3.86 33.32 6.31 55.90 True Context 5.13 35.26 8.08 54.14 Mixtral 8x7B No Context 3.34 40.41 5.65 60.20 Retriever 4.23 35.93 6.62 56.30 True Context 3.65 38.02 5.88 60.54
Zero shot QA.  The results for zero shot QA (see Table 5) show varying performance across different metrics. Notably, model size is not the sole determinant of performance. For instance, Llama-chat models, with fewer parameters, perform comparably to GPT-3.5 (e.g., 8.31 F1 for Llama-chat 70b vs. GPT-3.5 175B with 7.74). Some models, like Vicuna and Mistral, suggest a trade-off between precision and comprehensiveness. Models like Zephyr and Falcon, despite lower precision and F1 scores, have high recall and containment scores, indicating their ability to capture significant portions of the ground truth answers.
Few shot QA.  In the few shot learning setting [6], models are provided with a limited number of examples to guide their understanding of the task at hand. As demonstrated in Table 7, model performance improves with an increasing number of shots. This suggests that the examples help generate more accurate responses. For instance, the Llama-chat models with 7B, 13B, and 70B parameters show improvements as the number of shots increases from 0 to 3. However, the rate of improvement tends to plateau after 2 shots, indicating a diminishing return on the benefits of additional examples. The Mistral-Instruct and Mixtral models exhibit a similar trend.
RAG QA.  Retrieval-augmented generation (RAG) [26, 2] combines the strengths of pre-trained language models and information retrieval systems to generate responses in question-answering settings. RAG retrieves relevant documents from a large corpus when a question is posed, providing these documents as additional context to a language model to generate a response based on both the question and the retrieved documents.
In this experiment, we evaluate the efficiency of appending retrieved context to LLMs using Dense Passage Retriever (DPR) [24] for question answering. We test three settings for each model: without context, with the top retrieved passage from Wikipedia as context, and with the true context. The true context is determined by first retrieving 1,000 passages using DPR, then conducting a search within these passages, and selecting the highest-ranked passage containing the answer. If the answer is not found in any of the retrieved passages, a random passage is selected.
Table 7 shows the results. Model performance improves with provided context, with true context generally yielding the best outcomes. This suggests that relevant context enhances the accuracy and relevance of responses. However, performance varies across models and settings, indicating that RAG’s effectiveness depends on the specific model and the quality of the retrieved context.
Table 2: LLMs’ performance on Attribute-, Comparison-, and Counting-type questions.
Method Parameters Attribute-type Comparison-type Counting-type Precision Recall F1 Con Precision Recall F1 Con Precision Recall F1 Con Llama-chat 7B 6.55 60.70 9.52 96.86 5.30 58.05 8.83 63.20 0.34 10.50 0.66 80.24 Llama-chat 13B 6.17 61.26 10.67 96.77 5.22 59.53 8.75 63.30 0.16 5.86 0.30 86.13 Llama-chat 70B 8.21 65.85 13.71 97.63 7.06 61.41 11.25 64.76 0.43 16.42 0.84 48.99 Mistral-Instruct 7B 5.87 54.00 9.96 96.52 8.33 57.79 12.73 62.18 0.68 20.80 1.31 87.45 Mixtral 8*7B 4.04 75.80 7.21 97.25 5.18 63.50 8.86 72.10 0.60 27.44 1.17 90.10
Table 3: LLMs’ performance on Entity, Event, and Time questions.
Method Parameters Entity Event Time Precision Recall F1 Con Precision Recall F1 Con Precision Recall F1 Con Llama-chat 7B 4.26 40.86 6.95 75.80 4.75 53.25 7.87 82.00 3.38 41.86 5.84 76.09 Llama-chat 13B 3.79 40.70 6.52 80.95 4.96 49.72 8.31 81.80 3.14 38.27 5.45 81.28 Llama-chat 70B 5.07 47.55 8.45 82.31 7.23 60.18 11.28 85.10 4.50 53.50 7.67 84.02 Mistral-Instruct 7B 3.91 35.43 6.53 79.20 7.49 62.28 11.64 84.60 4.31 39.85 6.95 81.55 Mixtral 8*7B 2.82 50.95 4.85 84.76 4.91 60.06 8.29 86.60 2.62 54.78 4.77 85.85
Effectiveness of LLMs in Different QA Settings.  We evaluated the LLMs across different question types, including attribute, comparison, counting, entity, event, and time questions. As seen in Table 2 and 3, the models demonstrate varying performance across different question types, with attribute- and comparison-type questions proving easier to answer compared to counting-type questions. Specifically, the models perform better in answering questions related to events than entities, and they show the poorest performance with time-related questions.
7 Dataset Use
LLM Evaluation and Training. ComplexTempQA can be used for evaluating LLMs, as shown in our preliminary study in Section 6. Furthermore, it can be leveraged for fine-tuning Language Models, for prompt-based engineering, and for assessing the LLM’s capabilities, particularly in the area of temporal question answering [45].
Developing and evaluating truthful LLMs that can deal with hallucinations are a crucial research topic [35]. Our dataset allows to evaluate LLMs in this respect (concerning truthfulness) to an unprecedented degree, both regarding the diversity of questions (see our taxonomy) as well as the extent. Furthermore, it offers utility in the detection of hallucinations in LLMs, aiding in the evaluation and enhancement of models’ answer-generation capabilities. ComplexTempQA supports also the development of time-based datasets [9] due to its detailed temporal metadata, broadening the scope of temporal question-answering research.
Online Adaptation of LLMs. The development of online adaptation and continual training approaches to LLMs [17, 40] could also benefit from a large scale dataset like ComplexTempQA due to its temporal metadata which assigns the focus time scope to each question. On average, about 280k questions are associated with each year within ComplexTempQA’s time frame making it possible to experiment with the LLMs’ adaptation techniques to time. We note that ArchivalQA [47], the largest dataset used in [17, 40], has only about 22k questions on average per year.
RAG Systems. ComplexTempQA can be used to train and test Open Domain Question Answering (ODQA) models using (historical) news article archives, such as the New York Times (NYT) Annotated Archive [36], which contains 1.8 million articles published from 1987 to 2007 and has been widely used for temporal IR and related research [47, 5, 48]. Due to the NYT’s international coverage and the large number of US entities and events that ComplexTempQA focuses on, most questions should be answerable using the archive. In this regard, ComplexTempQA could serve as a complementary resource for answering temporal questions from the NYT or from other diachronic news corpora that align with ComplexTempQA’s time frame.
KGQA Systems. Built from Wikidata and Wikipedia, ComplexTempQA can be used for Question Answering over Knowledge Graphs [44, 37]. Integrating with large knowledge graphs such as Wikidata (cross-domain) and SemOpenAlex (covering publications) [12], both containing billions of facts, enhances the utility of QA models for exploring complex relationships between entities over time and evaluating complex question answering over multiple hops [15].
8 Conclusions
In this paper, we presented ComplexTempQA, a large dataset with over 100 million temporal question-answer pairs. ComplexTempQA significantly outperforms existing benchmarks in scope and coverage. Built on Wikipedia and Wikidata, the dataset covers questions from over three decades and offers a unique range of topics. We further introduced a novel taxonomy that categorizes the questions into attributes, comparisons, and count questions, with each question focusing on events, entities, and time periods. Another feature of ComplexTempQA is the high complexity of the questions, which requires skills such as across-temporal comparisons, temporal aggregation and multi-hop inference with temporal event ordering and entity recognition. Each question is accompanied by detailed metadata, including specific temporal domains, which enabled comprehensive evaluation and improvement of the temporal reasoning capabilities of large language models (LLMs).
Additionally, we evaluated various LLMs on our dataset, obtaining insights into their performance on temporal questions. Overall, ComplexTempQA serves as a valuable resource for developing sophisticated AI models and for advancing research in question answering, information retrieval, and language understanding. It is also useful, for example, for testing LLMs’ temporal factual knowledge, identifying the hallucinations they produce, and assessing their reasoning ability over time.
In the future we want to focus on extending datasets to other time periods and other cmplex question types, as well as consider insights from collective memory related studies [13, 19] to better select events and time periods for question generation.
Appendix A Example Questions and Templates
In our question generation process, we employed a range of templates, each tailored to different contexts and requirements. Below, we present several question templates showcasing the diversity of our approach:
What was [ATTR] of [ENTITY]?
- What was the genre of the movie Border?
What was [ATTR] of [EVENT]?
- What was the number of injured people of the aviation accident in 1988 which resulted in 143 deaths?
Who was [ENTITY] [YEAR]?
- Who was the founder of the Hess Corporation?
How many times did [PERSON] win [ENTITY] [TIME]?
- How many times did Robert Richardson win Academy Award for Best Cinematography between 1987 and 2007?
When did [PERSON] win [ENTITY]?
- When did Tom Hanks win the Academy Awards for Best Actor?
Who won [ENTITY] [YEAR]?
- Who won the the Academy Awards for Best Actor in 1999?
When was [PERSON] [ENTITY]?
- When was Girija Prasad Koirala President of Nepal?
Who was [ENTITY] [SIGNAL] [PERSON]?
- Who was Governor of California before Arnold Schwarzenegger?
Who was the winner of [ENTITY]?
- Who was the winner of the UEFA Champions League Final in 1999?
How often did [EVENT] happen [TIME]?
- How often did an aviation accident with more than 54 participants in Argentina happen in 1999?
Was [PERSON] [ENTITY] [YEAR]?
- Was Gary Locke Governor of Washington after Mike Lowry?
Who won [ENTITY] [YEAR]?
- Who won the Academy Award for Best Live Action Short Film in 1987?
In which country did the [EVENT] happen?
- In which country did the final of the FIFA World Cup in 1990 happen?
Did the [EVENT] happen in [COUNTRY]?
- Did the train wreck in 2000 that resulted in 155 deaths happen in Austria?
When did the [EVENT] happen?
- When did the Fall of the Berlin wall happen?
Did [EVENT1] happen [SIGNAL] [EVENT2]?
- Did the Gulf War happen after the Hurrican Hugo?
Did [EVENT1] or [EVENT2] happen first?
- Did Helios Airways Flight 522 accident or 27th G8 summit happen first?
Which one happened first, [EVENT1], [EVENT2], or [EVENT3]?
- Which one happened first, 69th Academy Awards, USS Cole bombing or Daegu subway fire?
What was the [ATTR] of [Entity]?
- What was the gender of Elton John?
Who was the [ATTR] of [EVENT]?
- What was the number of injured people of the car bombing in 1998 which resulted in 29 deaths?
Did [EVENT1] have a [COMPARE] [ATTR] than [EVENT2]?
- Did the car bombing in 1993 that resulted in 6 deaths have a higher number of death than the train wreck in 1989 that resulted in 645 deaths?
When [V1] [V2] [ENTITY]?
- When was the publication date of the movie where the actors Lou Diamond Phillips and Esai Morales played?
Did [V] [ENTITY1] [SIGNAL] [V] [ENTITY2]?
- Did the publication date of the movie where the actors Lou Diamond Phillips and Esai Morales played happened before the publication date of the movie Resident Evil: Extinction?
[V] [ENTITY1] or [ENTITY2] first?
- Was Modibo Keita, Ahmed Mohamed ag Hamani or Ousmane Issoufi Maïga Prime Minister of Mali first?
Which one [V] [ATTR] first, [ENTITY1] or [ENTITY2]?
- Which one had publication date first, the movie Alien vs. Predator or the movie where the actors Lou Diamond Phillips and Esai Morales played?
What was [V] of [ENTITY]?
- What was the country of origin of the movie Mudhalvan?
Did the [ATTR] of [ENTITY1] happen [COMPARE] the [ATTR] of [ENTITY2]?
- Did the publication date of the movie where the actors Lou Diamond Phillips and Esai Morales played happened before the publication date of the movie Resident Evil: Extinction?
Did [ENTITY1] have a [COMPARE] [ATTR] than [ENTITY2]?
- Did the the movie Jason Goes to Hell: The Final Friday had a higher duration than the the movie The Wind That Shakes the Barley?
Did [ENTITY1] or [ENTITY2] have a [COMPARE] [ATTR]?
- Did the the movie Jin-Roh: The Wolf Brigade or the the movie Ratatouille have a higher duration?
In how many years did [ENTITY] [DO] in [GENRE] [CAT] [ADDITIONAL] [YEAR]?
- In how many years did Jesse Ventura appear in a movie with Arnold Schwarzenegger between 1987 and 2007 with a duration of more than 60 minutes?
In how many years did [Event] [YEAR] happened?
- In how many years did a aviation accident by Pacific Southwest Airlines with more than 38 deaths between 1987 and 2007 happened?
In how many years did [ENTITY1] act with [ENTITY2] in [YEAR] [ADDITIONAL]?
- In how many years did Ron Howard act in movies with a duration of more than 60 minutes between 1987 and 2007?
What was the [P2] of the [P1] of [ENTITY] in meters?
- What was the highest point of the country where Moscow theater hostage crisis happened in meters?
Comparing [P1], has [ENTITY1] [COMPARE] [P2] than [ENTITY2]?
- Comparing country, has the company Meta Platforms a lower highest point than the company Renault?
Comparing [P1] of [ENTITY1] and [ENTITY2], which one has a [COMPARE] [P2]?
- Comparing country of the company Google and the company Yandex, which one has a lower highest point?
How many times [V] [PERSON] [ENTITY] [TIME]?
- How many times won Andrea Arnold the Academy Award for Best Live Action Short Film between 1987 and 2008?
Was [PERSON] [ENTITY] [YEAR]?
- Was Arnold Schwarzenegger Governor of California in 2001?
Table 4: Example questions with their answers in parentheses. Time questions are integrated within either events or entities which are indicated in square brackets.
Question Type What was the number of participants in cycling at the 1996 Summer Olympics – women’s sprint? (14) EventAttr When was the fall of the Berlin Wall? (1989-11-09,1989-11-09) TimeAttr [Event] Did the mass shooting in 1987 which resulted in 7 deaths have a higher number of injured people than the aviation accident in 1992 in Vietnam which resulted in 30 deaths? (yes) EventComp Did Halabja chemical attack happen after John F. Kennedy Jr.’s plane crash? (no) TimeComp [Event] How often did an aviation accident with more than 18 injured people between 1987 and 2007 in Spain happen? (3) EventCount In how many years did an aviation accident with more than 54 participants between 1987 and 2007 by Birgenair happen? (1) TimeCount [Event] Who was the director of the movie where the actors Bob Hoskins and John Leguizamo played? (Rocky Morton, Annabel Jankel, Roland Joffé, Dean Semler) EntityAttr Who was Emperor of Japan in 1987? (Hirohito) TimeAttr [Entity] Did the architectural structure which is designed by the architect KlingStubbins in 1990 have a lower amount of floors above ground than the Perlan? (no) EntityComp Which one had the date of birth first, the person Albert Zafy or the person Xanana Gusmão? (Albert Zafy) TimeComp [Entity] In how many years did Quentin Tarantino act with George Clooney between 1987 and 2007 with a duration of more than 60 minutes? (2) EntityCount In how many years did Michael Crichton write books between 1987 and 2007? (10) TimeCount [Entity]
In our question generation process, we employ placeholders to convey specific elements within the question templates. The placeholder [SIGNAL] denotes temporal relationships, such as ”before” or ”after,” indicating the sequence of events or entities. [ATTR] stands for attributes like the number of injured people, providing context or additional information related to the events or entities.[YEAR] referring either to a specific year or between two years e.g. ”in 1987” or ”between 1987 and 2007”. [COMPARE] signifies comparison relationships, such as ”higher” or ”lower,” enabling the comparison of attributes or characteristics between events, entities, or their attributes. Then we have [V] to denote various verbs like ”was” or ”happened”. [P] stands for properties used to create multi hop questions. We have one more complex template involving with [GENRE], [CAT] as categories (e.g., movie) and [ADDITIONAL] specifies an optional duration limit for movies. Furthermore, we make a clear distinction between questions involving persons and those involving other entities, as the question structure may vary accordingly. Despite these differences, some question templates are versatile enough to accommodate both event and entity questions, ensuring flexibility and adaptability in our approach to question generation.
Figure 8: Schematic of question creation pipeline: Utilizing Wikipedia and Wikidata, events and entities are extracted from page views, filtered, queried via SPARQL, and stored for analysis. Diverse questions, including multi-hop, unnamed event and counting-types, are generated, improves the dataset balance and complexity for evaluating question-answering systems.
Appendix B Insights in Metadata
We provide the following metadata: [^4] For the Wikidata IDs we exclude the leading ’Q’ as well for the properties the leading ’P’
Entity in question: A list of Wikidata IDs of the question.
Entity in answer: A list of Wikidata IDs of the answer if it contains an entity.
Country in question: A list of country Wikidata IDs of the countries of the questioned entities.
Country in answer: A list of country Wikidata IDs of the countries of the entities in the answer.
Hop property: A list of Wikidata properties of the question if it contains a hop. If there are multiple hops, they are listed in the order of use.
Rating: A numerical rating where 0 is considered an easy question, and 1 is considered a harder question according to the rating as described in Section 3.
Is unnamed: A numerical which is 1 if the question contains an implicitly described event and 0 otherwise.
Type: The type based on the taxonomy given in Figure 3.
Time span: The time frame to which the question relates to. For example, for the entity questions, the time frame ranges from born/creation to death/destruction. The start is always the earlier date, and the end is the latter.
Below is an example for a question including all the metadata:
  • Question: What was the highest point of the country where the 1988 Summer Olympics happened, in meters?
  • Answer: [1950]
  • Entity in question: [8470]
  • Entity in answer: []
  • Country in question: [884]
  • Country in answer: []
  • Hop property: [17, 610]
  • Rating: 1
  • Is unnamed: 0
  • Time span: [1988-09-17, 1988-10-02]
B.1 Event and entity popularity
Figure 9: Page views of entities and events (red dotted line marks the division between hard and easy entities and events).
Figure 9 shows the plots of entity and event popularity as measured by page views with threshold lines indicated to denote easy and hard separations.
B.2 Insights into events and entities
(a) Top categories of events.
(b) Top categories of entities.
(c) Top events by country.
Figure 10: Insight into events and entities
A closer examination of the categories reveals, as depicted in Figure 10(a), that our top event is aviation accidents, owing to their frequent occurrence in Wikipedia year pages. Subsequent events predominantly revolve around human crises such as wars or natural disasters, indicating a prevalence of negative events in our dataset. Moreover, a majority of these events transpired in the USA or other English-speaking nations, which is attributed to the use of English Wikipedia pages. As for entities, as illustrated in Figure 10(b), movies and persons feature predominantly in our questions.
Appendix C Experiment Prompts
In all experiments, each LLM was prompted to function as a helpful assistant and deliver direct, concise answers. For models employing the Retrieval Augmented Generation (RAG) approach, the prompt included additional context from retrieved documents to enhance the response quality. This modification ensured that responses were informed by relevant background information.
For all the experiments involving LLMs, we used the following prompt:
You are a helpful assistant. Provide direct and concise answer to the following question.
Question: <question>.
In the case of RAG, the prompt is slightly modified to incorporate the additional context provided by the retrieved documents. The prompt used for RAG is:
You are a helpful assistant. Using the context, provide direct and concise answers to the following question.
Question: <question>.
Context: <context>.
Appendix D Dataset creation pipeline
Figure 8 presents the detailed question creation pipeline.